OpenStudy (smart8thgrader):

A soup can has a height of 4 inches and a diameter of 3.5 inches. Which of the following is the closest to the volume of the soup can? 22 cubic inches 154 cubic inches 44 cubic inches 38 cubic inches

OpenStudy (smart8thgrader):

I WILL GIVE A MEDAL!!!!!!!!!!!

OpenStudy (smart8thgrader):

@brittanydosey

OpenStudy (smart8thgrader):

@k_lynn

OpenStudy (k_lynn):

The volume formula of a cylinder is \[\Pi r ^{2} h\] However the radius times itself is the diameter, so this is the formula you would use. \[\Pi d h\] Plug in 3.14 for \(\pi\) and then 3.5 for the diameter and 4 for the height.

OpenStudy (smart8thgrader):

@stephaniehoran23

OpenStudy (k_lynn):

Once you plug those in, what does the equation look like?

OpenStudy (smart8thgrader):

wait i dont get this period!!!!!!!!

OpenStudy (k_lynn):

Which part don't you get?

OpenStudy (k_lynn):

I'll try to explain it better.

OpenStudy (smart8thgrader):

ill fan if you tell me the answer@k_lynn

OpenStudy (k_lynn):

I'm not allowed to just give you the answer. That would be against the Code of Conduct.

OpenStudy (alexandervonhumboldt2):

formula is pi*r*r*h h=4 r=1/2*3.5=1.75 now substitue 1.75 for r and 4 for h for pi=3.14

OpenStudy (alexandervonhumboldt2):

you get 3.14*1.75*1.75*4=use calculator to solve this if you are confused with decimalss

OpenStudy (smart8thgrader):

i got 21.98\

OpenStudy (alexandervonhumboldt2):

nope

OpenStudy (alexandervonhumboldt2):

3.14*1.75*1.75*4=???????

OpenStudy (alexandervonhumboldt2):

no direct answers

OpenStudy (smart8thgrader):

thx

OpenStudy (alexandervonhumboldt2):

np